OpenStudy (anonymous):

Another answer check - What is the range of the following function f(x) = sqrt 2x +3 with the domain -1, 0, 1, 2, 3, 4 ,5? - My answer in the comments (also better problem view).

OpenStudy (anonymous):

Function: \[f(x) = \sqrt{2x+3}\] My answer, range:\[2, 0, \sqrt{5}, \sqrt{7}, 3, \sqrt{12}, \sqrt{13}\]

OpenStudy (dtan5457):

Looks like all you have to do is plug in your domain (x) into your function, but x=4 y=sqrt 11, if I'm not mistaken..?

OpenStudy (dtan5457):

x=-1 y=1 x=0 y=sq rt 3

OpenStudy (dtan5457):

everything looks about right

OpenStudy (dtan5457):

If this question did not have a set domain that range of that function would be all real non-negative numbers

OpenStudy (anonymous):

Sorry, I was afk for a bit. Can you show the solving process?

OpenStudy (dtan5457):

Well I mean it's exactly what you did..plug the x value into the function. I just see that you plugged x=4, and x=-1 wrong

OpenStudy (anonymous):

So it's sqrt 1 for x=-1 and sqrt 11 for x=4?

OpenStudy (dtan5457):

yeah.

OpenStudy (anonymous):

okay, thanks :D